Starting at age 25, Tracy invests $1500 at the end of each year in an IRA with an APR of 6.5% compounded annually. How much is in Tracy’s retirement account when she retires at the age of 65? (Assume the 40th and last deposit generates no interest).
Answer by stanbon(75887)

Starting at age 25, Tracy invests $1500 at the end of each year in an IRA with an APR of 6.5% compounded annually. How much is in Tracy’s retirement account when she retires at the age of 65? (Assume the 40th and last deposit generates no interest).
-----
A(t)=P(1+(r/n))^(nt)
---
A(40) = 1500(1+(0.065/1))^(1*40)
A(40) = 1500(1.065)^40
A(40) = $18624.11
=========================
Cheers,
Stan H.
=========================